OpenStudy (anonymous):

how to convert 100 yards to meters? ( please show the process)

OpenStudy (jdoe0001):

how many cm's in 1 inch?

OpenStudy (anonymous):

2.54

OpenStudy (jdoe0001):

site is a bit laggy...

OpenStudy (jdoe0001):

\(\bf \left(\cfrac{1\cancel{ in }}{2.54\cancel{ cm }}\cdot \cfrac{100\cancel{ cm }}{1m}\cdot \cfrac{1yd}{36\cancel{ in }}\right)\cdot 100\\ \qquad \qquad \uparrow \\ \textit{how many yards in 1meter} \)

OpenStudy (anonymous):

I dont get it

OpenStudy (jdoe0001):

notice the measures equivalents 1in to 2.54cm 100cm in 1m 1yd to 36in then you cancell those units out.. and mulitply the fractions to get how many "yards in 1 meter" you got 100 yards.... thus multiply that much times 100

OpenStudy (jdoe0001):

you'd usually put those equivalents in a fraction like above and simply put one above the other.... doesn't quite matter which goes above or below so you arrange them in a way that cancels out so in this case I wanted so cancel out say "cm" so in one of the fractions is below and in another fraction is above so it cancels out

OpenStudy (jdoe0001):

the unit cancels out....but the value.... needs to be multiplied across, like in any fraction multiplication

OpenStudy (jdoe0001):

so in the end the only units left, are "yards" and "meters"

OpenStudy (anonymous):

now i think i get it thanks

OpenStudy (jdoe0001):

\(\bf \left(\cfrac{1\cancel{ in }}{2.54\cancel{ cm }}\cdot \cfrac{100\cancel{ cm }}{1m}\cdot \cfrac{1yd}{36\cancel{ in }}\right)\cdot 100\\ \qquad \qquad \uparrow \\ \textit{how many yards in 1meter} \\ \quad \\ \implies \left(\cfrac{1}{2.54}\cdot \cfrac{100}{1}\cdot \cfrac{1}{36}\right)\cdot 100\implies \left(\cfrac{1\cdot 100\cdot 1}{2.54\cdot 1\cdot 36}\right)\cdot 100\)

OpenStudy (jdoe0001):

yw